### v1

# Future Trends in the Design of Memetic Algorithms: the Case of the Linear Ordering Problem

Abstract
The way heuristic optimizers are designed has evolved over the decades, as computing power has increased. Initially, trajectory metaheuristics used to shape the state of the art in many problems, whereas today, population-based mechanisms tend to be more effective. Such has been the case for the Linear Ordering Problem (LOP), a field in which strategies such as Iterated Local Search and Variable Neighborhood Search led the way during the 1990s, but which have now been surpassed by evolutionary and memetic schemes. This paper focuses on understanding how the design of LOP optimizers will change in the future, as computing power continues to increase, yielding two main contributions. On the one hand, a metaheuristic was designed that is capable of effectively exploiting a large amount of computational resources, specifically, computing power equivalent to what a recent core can output during runs lasting over four months. Our analysis of this aspect relied on parallelization, and allowed us to conclude that as the power of the computational resources increases, it will be necessary to boost the capacities of the intensification methods applied in the memetic algorithms to keep the population from stagnating. And on the other, the best-known results for today’s most challenging set of instances (xLOLIB2) were significantly outperformed. Instances with sizes ranging from 300 to 1000 were analyzed, and new bounds were established that provide a frame of reference for future research.
keywords:
Linear Ordering Problem , Memetic Algorithms , High-performance Capabilities , Design of Optimizers1 Introduction
An effective decision-making process usually involves solving highly complex optimization problems [1]. The most general classification of optimization techniques distinguishes between exact [2] and approximate [3] optimization. Given the complexity of many real-life optimization problems, many of which are NP-hard, exact optimization methods are usually not applicable within reasonable timeframes [4]. Therefore, approximate methods that strike a balance between the quality of the solution achieved and the computational cost have gained considerable popularity [5]. In this area, metaheuristics offer the possibility of increasing the generality of the developments, thus reducing the effort needed to adapt optimizers to new optimization problems.
One highly complex problem that has benefited greatly from the use of metaheuristics is the Linear Ordering Problem (LOP), which is an NP-hard combinatorial optimization problem with a large number of applications. The LOP was introduced in 1936 by Leontief [6] and formalized in 1958 by Chenery and Watanabe [7]. The simplest definition of it is based on the matrix triangulation problem. Given a matrix , the problem is to determine a simultaneous permutation of the rows and columns of such that the sum of the entries above the main diagonal is maximized [8]. As a result, the problem can be formulated as the search for a permutation that maximizes the equation:
Although the LOP originated in the field of economics, it has been applied in numerous areas, such as: determination of ancestry relationships in archaeology [9], straightline crossing minimization [10], and the development of probabilistic decision-making models [11]. Due to its broad applicability, the LOP has received considerable attention from the scientific community in recent decades [12, 13]. Moreover, since many of these applications involve the use of large matrices, highly complex test sets have been defined. In particular, the most challenging validation set today is the LOLIB library [14], the latest version of which includes instances up to size 1000.
Many diverse algorithms have been proposed to deal with the LOP [8]. This is because, in general, the design of metaheuristics is a complex undertaking, since most of them rely on stochastic processes that are difficult to model and analyze theoretically. Because of this, theoretical analyses often resort to simplifications, meaning they have a restricted scope and the design of optimizers cannot be based solely on theoretical advances [15]. This means that when designing new metaheuristics, researchers must draw on the practical experience gained from similar problems, since they do not have a specific guide of steps to follow [16]. In addition, the relative efficiency between metaheuristics depends largely on the computational capabilities available and the complexity of the instance to be solved, which accounts for the need for a wide variety of proposals [17]. All this explains why the leading metaheuristics designed a few decades ago are very different from the more successful metaheuristics designed in more recent times.
The LOP is a fairly clear example of how computational resources and the complexity of instances affect the design of optimizers. Thus, based on the historical review presented in [8], it is clear that effective optimizers have been evolving from ad-hoc heuristic methods, to trajectory metaheuristics, and on to the current population-based metaheuristics [18, 19].
This paper falls within the field of approximate optimization and aims to contribute to the advancement of the design and development of efficient and robust optimizers. It focuses on the case of the LOP and, in particular, seeks to answer the following research questions:
1.
Are the methods that generate the best solutions for the LOP able to do so in reasonable times, for example, under a day in the most challenging instances currently in use?
2.
Will today’s cutting-edge methods in the LOP continue to retain their prominence in the computing systems of the future, that is, after a significant increase in computing capabilities? Or, on the contrary, will they be unable to scale properly and require methodological changes?
3.
Have optimal solutions already been achieved for the current most challenging datasets, or, on the contrary, is there still room for improvement in the design of the solvers for the LOP?
In order to answer these questions at this time, it is necessary to make use of long-term executions and parallelizations to simulate the increased availability of computational resources. With this goal in mind, we conducted a long-term performance analysis of the current leading method for the LOP, the MA-EDM algorithm [19]. This analysis allowed us to identify some shortcomings in this method, and led us to conclude that, as is usually the case, once the computing power is increased, the algorithms have to be redesigned in order to leverage the additional computing power. In light of the weaknesses identified in the MA-EDM, a new optimizer called Memetic Algorithm with Explicit Diversity Management and Enhanced Intensification (MA-EDM) is proposed that applies an Iterated Local Search (ILS) as an improvement mechanism. As demonstrated in the experimental study, in high-performance environments, MA-EDM vastly outperforms the current best strategy. Moreover, by parallelizing MA-EDM, we were able to simulate scenarios with computing capabilities equivalent to those of a state-of-the-art core in executions lasting over 4 months. This experiment revealed that the best solutions found so far exhibit a significant gap with respect to the optimal ones, indicating that this set of instances has not yet been optimally solved, meaning, therefore, that more methodological advances are required. As an additional contribution, it should be noted that the new Best-Known Solutions (BKS) obtained for instances ranging in size from 300 to 1,000 will provide a frame of reference that will allow other researchers to better evaluate future advances.
The rest of the document is organized as follows. Section 2 discusses the design of approximate algorithms, with a special emphasis on Memetic Algorithms and their adaptations to the LOP. Section 3 focuses on describing the use of high-performance computing environments to achieve the objectives of this research. Section 4 justifies and presents a new algorithmic proposal for the LOP called Memetic Algorithm with Explicit Diversity Management and Enhanced Intensification (MA-EDM). The experimental validation is presented in Section 5. Finally, the last section discusses the conclusions of this research, as well as some potential future areas of work.
2 Future and Limitations of Memetic Algorithms for the LOP
Evolutionary Algorithms (EAs) are a class of metaheuristics that draw their inspiration from biological evolution [20, 21]. Experience has shown that EAs are very effective at solving difficult optimization problems in a wide variety of fields [22]. However, they typically become more effective when combined with trajectory techniques, or some other refinement mechanisms [23]. This combination of EAs with additional schemes is known as Memetic Algorithms (MAs) [24].
Multiple strategies have been designed to integrate EAs with other mechanisms. In first-generation MAs, enhancement mechanisms were used statically to refine the solutions found by genetic operators. Subsequently, a variety of more complex integration strategies were developed, ranging from schemes that consider a set of improvement mechanisms that are applied adaptively, to schemes that rely on the feedback provided by the improvement mechanisms to adapt internal parts of the EA, such as its fitness function or genetic operators. Despite the advances offered by recent MAs [25], their design, use and implementation is very complex due to the number of parameters and components they consider, as well as to the complex interactions that occur between them. In many problems, the first-generation MAs are still very effective [26, 27, 18], so both the simplest and the most complex methodologies continue to be investigated.
Considering the number of problems in which MAs are the leading strategy [21], and that this is the case with the LOP, this paper focuses on these types of strategies. In particular, in the case of the LOP, the MA that has yielded the current BKS is a first-generation algorithm [19], so this is the method that will be used to conduct the proposed analyses.
The rationale behind the development of MAs is that EAs are often appropriate for exploring different regions and identifying promising areas, but their intensification capabilities are more limited. Because of this, the idea was to integrate EAs with optimizers that had good intensification capabilities. In this sense, hill climbing local search is one of the simplest schemes to intensify a region of the search space. Since, for many problems, finding a local optimum is a computationally very expensive process, in such cases it is not feasible to use more complex schemes, which is why local search remains one of the most used methods in MAs; however, advances in hardware have also made it quite common for more complex trajectory metaheuristics, such as Tabu Search (TS) or Variable Neighborhood Search (VNS), to be considered, and even mathematical programming strategies that are usually even more computationally expensive.
In relation to the design of MAs, the problem of how often and how the improvement mechanism is used is a fundamental task that has been addressed in the literature in various ways. In the context of continuous optimization, the frequency and application point of local searches have been analyzed [28]. This analysis was expanded in [29] for combinatorial optimization problems.
Another crucial point is the definition of the neighborhood for the local search. In order to aid in the selection of the neighborhood, some heuristic procedures for analyzing the fitness landscape have been devised [24]. An in-depth study of neighborhood selection has been proposed, as well as a taxonomy that can be used to classify MAs, taking into account not only the way in which local search is applied, but also the interactions that arise between the different components [30]. It is also important to note that in addition to the design issues implicit to MAs, as with any population-based algorithm, there are several additional characteristics that play an important role in the success of the optimization process. In particular, proper diversity management plays a critical role [31, 32]. Diversity is a property of a group of individuals that quantifies how much they differ from one another. Although there are a variety of principles that can be followed when designing EAs, in the authors’ opinion, a population-based algorithm should ideally work in conditions of high diversity during the initial stages of optimization, and then progressively lose diversity in the vicinity of the global attraction basin. This position has been evaluated for a variety of combinatorial optimization problems. By following this design principle, algorithms have been developed that currently provide the BKS for different optimization problems, such as the Job Shop Scheduling Problem [33] and the Frequency Assignment Problem [34]. The ideal behavior discussed above hardly occurs in practice in EAs, since in most cases, the algorithm tends to converge prematurely towards a suboptimal solution, or to stagnate [32].
In the case of MAs, problems with diversity management appear frequently. For example, in the RHGA algorithm [25], one of the most effective strategies for the TSP, great care had to be taken in selecting the frequency of application of the LKH improvement strategy to avoid premature convergence. This type of problem arises especially in the case of high dimensionality problems [35], so as the instances grow, it is quite common to have to change some design decisions. Issues related to instances of high dimensionality are not unique to MAs [36]. This problem is known as the “curse of dimensionality”, a term proposed by Bellman in 1957 [37], and also appears in other types of optimizers and in machine learning. The curse of dimensionality occurs when an increase in search space size causes a significant loss of performance in optimization algorithms. The increase in the dimensionality or size of the instance causes the search space to grow exponentially in most problems, and as a result, the percentage of solutions evaluated decreases. However, this is not its only drawback, since it can, for example, cause premature convergence in subsets of variables [35]. Therefore, increasing the amount of computing power available (even exponentially) may not solve the challenges that arise when dealing with large instances. To address this problem, various techniques, such as reducing dimensionality and coevolution, have been proposed [38, 39]. This paper takes advantage of the demands associated with high dimensionality problems to fully leverage the metaheuristic schemes known to perform well with the LOP. To do this, we will work with instances of up to , that is, with a search space with approximately candidate solutions, and we will give the algorithm what it supposedly needs: more computing power. It is not at all foolhardy to think that in the near future, computing capacity will advance significantly, which will make more computing power available to algorithms. However, what we do not know is whether today’s methods will be able to continue to perform well when these resources are made available to them. The experiments conducted provide an answer to this question.
3 Using High Performance Computing Environments
High performance computing (HPC) laboratories consist of a set of multi-core machines that can communicate over high-bandwidth low-latency networks. These laboratories are typically used to reduce the time required to solve highly complex problems. Since the programming and bottlenecks of parallel algorithms differ from those of sequential algorithms, it is quite common for parallel algorithms to be developed that are significantly different from the direct parallelizations of the best sequential algorithms. Thus, in the case of Memetic Algorithms and population-based optimizers in general, various techniques have emerged to take advantage of this type of resource [40], such as master-worker parallelizations, island models, coevolution and cellular models. Among them, the only strategy that maintains the panmictic population-based model of the sequential algorithm is the master-worker model, while the others include changes in the evolutionary process that facilitate the parallelization and treatment of possible bottlenecks.
The objective of using HPC laboratories in this research was not to take advantage of parallel computing resources to obtain high-quality results quickly, but to understand how the effectiveness of one of the leading algorithms in the area of LOP is affected by the increase in computing capabilities, as well as to propose modifications that allow the increase in computing power to be better utilized. Thus, the purpose of this work is not to innovate in parallelization models or find the best parallel models for the LOP, but to generate new knowledge in relation to the optimization of the LOP so as to understand what type of research might be most useful to continue advancing in this field.
As mentioned, dimensionality is one of the characteristics that correlates highly with the difficulty of solving a problem. The case of the LOP is no exception, and over the years, new sets of instances have had to be created that pose a challenge to the best existing algorithms [41]. Although various characteristics have been identified in the instances and the search landscape that affect the difficulty of the solution, the size of the matrix is the most important and easily modifiable [42]. The eXtended LOLIB library (xLOLIB) [42] was proposed for the purpose of creating instances that are complex and exhibit characteristics similar to the matrices that arise in practical problems. Over the years, it has been updated with new instances, and the xLOLIB2 subgroup currently contains instances ranging from 300 to 1000. In recent years, the Best-Known Solutions (BKS) for the largest instances have continued to be updated [43, 19], which is an indication that there is still room for improvement for LOP optimizers when dealing with instances of these sizes. Specifically, MA-EDM [19] is the algorithm that has generated all the BKS for this subgroup of instances, so it is the algorithm that will be used in the analyses.
Considering that one of the purposes of this research is to understand how the MA-EDM and its extensions will behave when using higher performance processors, models should be used that keep the evolutionary process intact. Because of this, the synchronous master-worker model was selected. This scheme is standard in the parallelization of population-based algorithms, and can be programmed using what is commonly known as an embarrassingly parallel model, since the only extension needed is the redistribution of a set of independent tasks among a set of processes that are exclusively dedicated to executing those tasks. In order to facilitate the reproducibility of this research, this section details how to perform the parallelization.
The MA-EDM sequential algorithm is a standard first-generation MA. Its general behavior is described in Algorithm 1. A detailed description of each of its components is available at [19]. Each iteration operates with a population of individuals and consists of a binary tournament-based parent selection phase, a child-reproduction phase that considers the Cycle Crossover (CX), an improvement phase that applies a local search, and a replacement phase that uses the Best-Non Penalized strategy (BNP). Of all the above steps, the local search is the most computationally demanding process, since the neighborhood is considered based on shifts [44], so for instances of size , the step of ensuring that a local optimum has been found using incremental evaluation has a computational cost of — and this is only one part of the computation that is required in the local search — while the rest of the components depend linearly or sublinearly on . Therefore, with large instances like the ones considered in this research, it is natural for the improvement process to be the task that is distributed in the parallelization. Finally, and as mentioned, this distribution is done synchronously so as not to change the evolutionary process of the sequential scheme; that is, once a local search distribution point (or another improvement procedure) is reached, the master process halts until all the local searches are complete.
The flowchart shown in Figure 1 illustrates the operation of the parallel version of MA-EDM, an algorithm that is called Parallel MA-EDM (PMA-EDM). The diagram shows an example with a population of size 10, in which one master process (bottom) and four worker processes (top) are executed. The population is represented with larger squares, while the offspring is represented with smaller squares. The rectangular blocks in the region of the master process are executed sequentially, while the procedures performed in parallel are represented using two horizontal lines, and extend to the region corresponding to the workers. There are two stages in which the intensification procedure is applied to a set of individuals: after creating the initial population, and after creating the offspring. Those two blocks are the ones that are executed in parallel. By using the Message Passing Interface (MPI) [45], the individuals are distributed so that the worker processes carry out the intensification process and output the results. This distribution is carried out dynamically, meaning we wait for a process to finish carrying out the current intensification before sending it a new individual, continuing with this strategy as long as there are individuals remaining to be intensified. In the parallelized blocks, i.e., in the “Intensify”, the individuals that have already been sent to the work processes and were intensified are shown in green, those in which the intensification process is taking place are gray, and those that have not yet been intensified are red. The process will continue until the intensification has been performed on all the individuals, at which point the master process will continue with the next sequential blocks. The advantage of doing it dynamically is that it avoids imbalance problems when the tasks are heterogeneous. Such is the case with the local search, since the number of iterations to converge depends on the initial solution. Given that in the case of the LOP, the degree of heterogeneity is not high, in this research we selected population sizes that are multiples of the number of workers, which is an important consideration to enhance the efficiency. Finally, it is important to note that this way of parallelizing is independent of the components selected to carry out the intensification process, which is why the flowchart is presented in a general way. The parallelization shown was used to parallelize an extension of the MA-EDM algorithm presented in the next section.
4 Adaptations of MA-EDM to Improve its Long-Term Effectiveness
As computing power increases and the number of instances grows, it is quite common for the effectiveness of algorithms to be affected [17]. Different options to deal with the increase in computing power, comparing the alternatives for expanding the number of generations or the size of the population have been analized [46]. Based on its analyses, both theoretical and practical reasons are given in favor of increasing the size of the population and keeping the number of generations from growing excessively. There are currently attempts to develop effective population-based algorithms with large populations, since this would greatly facilitate the efficient use of computational resources, especially in parallel computing environments [47]. However, the use of large populations is still in the minority. Moreover, we know that by introducing mechanisms that delay premature convergence, the number of useful generations can be increased [32]. For example, some optimizers that use Best Non-Penalized (BNP) strategy — such as MA-EDM — have been able to successfully leverage several thousand generations [34]. However, even with this type of mechanism, there are limits on the number of useful generations, so with large computing capacities, some adaptations are still necessary.
The options to take advantage of the increased computing power to manage larger populations or handle more generations do not entail any type of change in the algorithm. A different option that will also be evaluated in this research to avoid the evolution of such a large number of generations is to introduce computationally effective (and costly) strategies in some of the components that are used in each iteration of the algorithm. In particular, for the improvement process, instead of applying a local search whose criterion is based on reaching a local optimum, a trajectory metaheuristic can be used to set a time-based stopping criterion. This option is not promising at present for the LOP with large instances because with the current amount of power, it would require setting a very short time or having very few generations, but as the computing power increases, this option becomes viable. This methodology makes it possible to control the number of generations without having to resort to increasing the size of the population by an inordinate amount.
Before being displaced by Memetic Algorithms, the Iterated Local Search (ILS) was one of the most successful strategies for the LOP [42], and to this day it continues to offer advantages in some variants of the LOP [48], so we decided to resume this research and include it within MA-EDM. Thus, we define the Memetic Algorithm with Explicit Management of Diversity and Enhanced Intensification (MA-EDM) with the novelty that in the intensification step, the local search is replaced with the ILS.
In order to properly design an ILS strategy, we must select the perturbation and local search components. Given the efficiency and effectiveness of the local search based on shifts, this was left unchanged. In the case of the perturbation strategy, we analyzed several operators that have been successful in combinatorial optimization [49] and that exhibit different perturbation strengths. In particular, we tried strategies widely used in the field of ILS: Scramble, Inversion, Insertion and Swap. In the case of the swap strategy, we considered the version that incorporates the parameter to establish the number of 2-position swaps that are made [50]. This parameter is used to model the strength of the perturbation. The preliminary analyses allowed us to determine that the method based on performing multiple swaps was far superior. In particular, was enough to escape local optima, so the rest of the paper considers this operator exclusively. In fact, this operator had already shown its superiority in [48], but since in this research, the ILS is integrated into the MA, we opted to obtain an independent confirmation.
To extend MA-EDM (Algorithm 1) and generate MA-EDM, we need only change the intensification steps (lines 2 and 7) by replacing the local search with the ILS. The Algorithm 2 details each step of the ILS. Note that it is the standard variant of ILS in which each solution is intensified with a local search after being perturbed, and it is accepted if it is at least as good as the best solution found. In relation to the parallel version of this algorithm, it follows exactly the same diagram presented above (see Figure 1). The only difference is that the workers execute the ILS instead of the local search. The parallel version is called PMA-EDM.
5 Experimental Validation
The main objective of the experimental validation carried out in this research is to analyze how well the MA-EDM algorithm can be scaled up with an increase in computing power, as well as to understand how its parameterization or its components can be changed to contend with this situation. As a secondary objective, we intend to apply the algorithms developed to the most challenging instances of the LOP by simulating scenarios with a large amount of computation, in order to understand the room for improvement present compared to the current BKS [19], as well as to establish new levels that can be used in the future to validate new advances in this area. The results are presented in two sections, each one focused on each of the objectives listed above.
To account for the stochastic behavior of the algorithms analyzed, the runs were repeated 30 times in all the experiments, and the following statistical analyses were performed to compare the results [51]. First, the Kruskal Wallis test was applied as an omnibus test [52]. If there were any significant differences, the Mann Whitney with Hommel’s correction was used to identify the pair-wise differences. Additionally, with the aim of visualizing the results, the Nemenyi critical difference plot was used [53], with a significance level equal to 0.05. The experimental validation was carried out in a cluster whose nodes have two Intel Xeon E5 v4 family processors and 32 GB of RAM.
5.1 Relationship between the scalability and components of MA-EDM
In order to analyze the scalability of MA-EDM in anticipation of a future increase in processor computing power, the increase is simulated in two ways: by increasing the stopping criterion, or through parallelization. The former yields an exact analysis of what will happen with an increase in computing power, allowing for a more limited range of increase, while the parallelization simulates a much larger increase in resources, although when there is a loss of efficiency, the value of the increase in computing power is approximate. It should be noted that the parallelization used keeps intact the evolutionary process of the MA that is parallelized.
First, in order to validate the benefit of increasing the number of generations, the performance of the MA-EDM and MA-EDM sequential implementations was analyzed using different stopping criteria (6, 24 and 192 hours). In order to have instances with different complexities, four instances of the xLOLIB2 package were chosen at random, one for each size available: 300, 500, 750 and 1000. The instances used were: N-t65i11xx_300, N-be75eec_500, N-stabu75_750 and N-t70b11xx_1000. Figure 2 shows the boxplots of the results obtained in each case. The only parameter of the MA-EDM algorithm is the population size (). Taking into account the results in [19], was used. In the case of MA-EDM, the same population size was used and the stopping criterion for ILS was set at 3.6 seconds. The preliminary analyses showed that after this run time, the improvements offered by ILS are quite limited. With this value, each generation lasts about 12 minutes, so in the longest runs, about generations evolve. In all the instances, it is evident that in runs with more limited stopping criteria (6 and 24 hours), the MA-EDM algorithm is superior. This is especially clear in the 6-hour run, a case in which there are significant differences in favor of MA-EDM in the 4 instances. However, as the computing power increases, the improvements achieved by the MA-EDM algorithm are quite limited, and it is far surpassed by the MA-EDM algorithm. This improvement is most noticeable for larger instances. In fact, in the instances with , and , the statistical tests mentioned earlier confirm the superiority of MA-EDM.
The advantages offered by MA-EDM in instances of size are more limited. The increase in computing power does not yield an increase in the fitness of the best solutions achieved, but it does produce a noticeable decrease in the standard deviation (both in MA-EDM and MA-EDM), so it is very possible that solutions very close to optimal are being achieved in instances of this size. In the cases of , and , the improvement provided by MA-EDM is broad, with even the worst result achieved by MA-EDM being better than the average achieved by MA-EDM. In the case of the instance of size , MA-EDM and MA-EDM evolved on average and generations, respectively. The results show that as the complexity of the instances grows, expanding the number of generations is not the best way to take advantage of the increased computational resources. In fact, in the case of the instance, despite the noticeable margin for improvement still left for MA-EDM, between the results obtained in 24 hours and 192 hours, i.e., when multiplying the computation amount by 8, there were no significant differences.
Taking into account the analyses carried out in [46], it is natural to consider MAs with large populations as an alternative to leverage the increase in computing power. Because of this, MA-EDM was run by setting the stopping criterion to 192 hours and using larger population sizes. In particular, 6 values were tested between and . The boxplots of the results obtained for the instances N-stabu75_750 and N-t70b11xx_1000 are shown in Figure 3. The results obtained by MA-EDM are also included. The statistical tests carried out confirm that in both instances, there were significant differences between the results achieved by MA-EDM and those obtained by MA-EDM, regardless of the population size used. As expected, there was a decrease in the number of generations. As an example, MA-EDM with averaged generations for instances of size , while increasing to reduced the number of generations to . However, this was not reflected in a so significant improvement in the quality of the solutions. We conclude that in this case, decreasing the number of generations through population growth is not enough to take advantage of the computing resources; instead, more drastic changes must be made by modifying some of its internal components, as is done in MA-EDM.
5.2 Analysis of performance in the very long-term
The above experiment showed that MA-EDM provides significant improvements in instances of size or larger. This second experiment seeks to establish new BKS and analyze whether MA-EDM scales appropriately when considering larger computation quantities than those used in the previous experiment. Since the previous experiment involved optimizations lasting up to 8 days, it is not feasible to perform this new experiment using sequential executions. Instead, PMA-EDM with 32 worker processes was used. One modification was made to the population size, consisting of establishing in order for the population size to be a multiple of the number of workers. Figure 4 shows the boxplots obtained by PMA-EDM when executed setting the stopping time between 6 hours and 26 hours for the 750 and 1000 size instances that were used in the previous experiment. Also shown are the results obtained by MA-EDM at 192 hours. The results for the 300 and 500 size instances are shown in the Supplementary Material. Note that the computation hours used in the parallel cases range from 192 to 832 hours. As expected, the 6-hour parallel results were worse than the 192-hour sequential results due to the loss of performance inherent in the parallel models. However, when using 10-hour runs, the median obtained by PMA-EDM was higher than that of MA-EDM. Therefore, the acceleration is estimated to be between and . When considering the mean value, the efficiency is equal to 0.8, and therefore each hour of parallel execution is estimated to correspond to hours of sequential execution. Thus, this experiment simulates situations with current computing capabilities between 153 hours and 665 hours, i.e., practically one month of execution.
In the case of the instance of size , the results continue to improve steadily, even on the longest runs. The improvement in quality obtained compared to the previous BKS is immense, thus confirming that in instances of size with the usual execution duration — at most a few days — there is considerable room for improvement in the LOP. Figure 5 shows the plots of the critical difference for these runs for the and instances. These same figures for the case of 500 are shown in the Supplementary Material of this research. In the case of 300, it is not possible to construct this graph since the same result was obtained for all the executions carried out with both algorithms. In both instances, we see that the sequential run at 192 hours ranked between the parallel runs at 6 and 10 hours, confirming the previous findings in terms of the estimated efficiency. In the 750 instance, 22-hour parallel runs were needed to significantly outperform the sequential results. When a method generates near-optimal solutions, the resources allocated have to be greatly increased to achieve significant improvements, so this, coupled with the fact that the 22-hour run ranked similar to the 26-hour run, suggests that in this instance, very high quality solutions are being reached or that the method developed is no longer capable of yielding additional improvements. In the case of the instance, the sequential results at 192 hours could be significantly outperformed by all the parallel runs starting at 14 hours. In addition, the best average ranking was obtained at 26 hours, which is an indication that there is still room for improvement.
| Instance | PMA-EDM | Previous BKS | ||
| Mean | Best | Worst | ||
| N-t65i11xx_300 | 864234341.0 | 864234341 | 864234341 | 864223106 |
| N-be75eec_500 | 33489159.76 | 33489269 | 33488116 | 33464804 |
| N-t59f11xx_750 | 29202835.91 | 29208868 | 29196802 | 29192928 |
| N-t59i11xx_1000 | 3493537276.58 | 3493926328 | 3493030762 | 3491553089 |
| N-t65l11xx_750 | 4948911.75 | 4949817 | 4948170 | 4944869 |
| N-t65w11xx_1000 | 72107414598.25 | 72127664540 | 72088824319 | 72045429648 |
| N-t75u11xx_750 | 17369141424.0 | 17372345327 | 17364459304 | 17358027725 |
| N-t75u11xx_1000 | 29732758325.58 | 29735696136 | 29727937321 | 29713770054 |
| N-tiw56n72_750 | 77341956.33 | 77356607 | 77324903 | 77300418 |
| N-tiw56r54_1000 | 29320651.75 | 29324010 | 29317329 | 29300654 |
| N-stabu75_750 | 91121628.5 | 91144824 | 91105532 | 91056055 |
| N-t70b11xx_1000 | 314863785.0 | 314977289 | 314792889 | 314603886 |
| N-tiw56r66_750 | 35482250.91 | 35484408 | 35477649 | 35461293 |
| N-tiw56r67_1000 | 66711779.25 | 66724580 | 66701153 | 66667742 |
In order to establish new levels that can be used in future research, PMA-EDM was executed for a set of 10 additional instances, with the stopping criterion set at 26 hours. Table 1 shows the mean, best and worst values achieved, as well as the BKS prior to this research [19]. This table also details the results obtained in the instances discussed previously. Note that in every case, the previous BKS was surpassed. In fact, even the worst result is better than the previous BKS. Also worth noting is that in the and instances, the margin for improvement is relatively small, but grows to several million in the -size instances.
| Instance | PMA-EDM | BKS | ||
| Mean | Best | Worst | ||
| N-stabu75_750 | 91127572.35 | 91150565 | 91105008 | 91056055 |
| N-t75e11xx_750 | 823367015.6 | 823481759 | 823055676 | 822729527 |
| N-usa79_750 | 158342573.25 | 158388557 | 158286792 | 158187746 |
| N-t70b11xx_1000 | 314906869.75 | 314989031 | 314817826 | 314603886 |
| N-t59n11xx_1000 | 9624726.95 | 9626140 | 9623842 | 9617008 |
| N-t70k11xx_1000 | 28550270180.0 | 28558080100 | 28544793200 | 28520983800 |
As a final experiment, and in order to set new BKS that are even more challenging to achieve, PMA-EDM was executed with 6 instances and 32 worker processes, with the stopping criterion set at 120 hours. These results are shown in Table 2. In light of the ratios calculated earlier, these runs correspond to sequential optimizations lasting over 4 months. Two of the instances are common to those used in the previous experiments, and the results continued to improve drastically in the case of , thus confirming that the current methods still have considerable room for improvement with instances of this size, and that in order to achieve solutions of this level of quality, truly long executions would be necessary.
6 Conclusions
The design of metaheuristics has evolved over time due to the emergence of new ideas to perform optimization processes, as well as the continuous increase in computing capabilities. This paper focuses on understanding the impact that the future increase in the availability of computational resources will have on Memetic Algorithms (MAs) in the particular case of the Linear Ordering Problem (LOP). The use of parallelizations that keep the evolutionary model intact, as well as of long-term executions, made it possible to confirm that the MA that is used today to obtain the Best-Known Solutions (BKS) in the most challenging instances of the LOP (MA-EDM) is not able to adequately leverage the increase in computational resources on a large scale. Two of the most common techniques used to take advantage of the increase in computational resources are to expand the size of the population or the number of generations. However, with a high increase in computational resources, neither of these two techniques was successful in the case of the LOP. A new heuristic optimizer (MA-EDM) was proposed based on integrating one of the most successful schemes of the 1990s (Iterated Local Search) with MA-EDM, so as to avoid an excessive increase in the population size or in the number of generations, and thus increase the intensification potential. This technique was able to successfully take advantage of amounts of computational resources similar to what a modern core can execute in more than 4 months, and yielded new solutions that significantly outperform the current BKS in the 18 instances that were used in this research. These achievements confirm that as computational resources continue to increase, it will be necessary to adapt the current optimizers, and in particular it shows how promising it is to integrate MAs with techniques with great intensifying power. Accordingly, we believe that in the future there will be a significant boom in MAs that will include intensification procedures with a high computational cost. In fact, in recent years, the term matheuristics has become popular to refer to heuristics that include mathematical programming techniques in some of their components. These types of techniques were not viable in the past due to their computational cost, but the number of areas in which this type of integration has been successful is growing. Also of note is the fact that the generation of BKS has been possible without incorporating initialization mechanisms or genetic operators that bias the search towards certain types of regions that might seem more promising. These types of ad-hoc components have been successful in the past. However, with the increase in computational resources, more regions can be explored, so the inclusion of some type of bias in the search, specially in the initial phases, can be counterproductive. Therefore, we predict that in the future, operators with a high degree of randomness - such as those used in this paper - will become more effective.
In relation to the LOP, this research also shows that the current instances are still very challenging, and that there is still much room for improvement. There is a need to integrate more methodological advances, mainly to be able to successfully deal with instances with matrices of size or greater, since in many fields it is not feasible to execute algorithms that are computationally equivalent to running a modern core for 4 months. In addition, given that even with the longest executions carried out for this work, significant improvements continue to be obtained — especially in the case of size — it is very likely that the optimal solutions can still be improved with respect to what is reported in this paper.
There are a large number of problems in which MAs very similar to those analyzed in this research make up the state-of-the-art methods. In the future, it would be interesting to explore whether the same drawbacks in terms of scalability appear in these areas, and if the same types of modifications help alleviate these problems. In relation to the LOP, it seems promising to further increase the capacities of the intensification process, so ideas related to the use of dynamic programming and branch-and-bound within the evolutionary process will be explored.